OpenStudy (anonymous):

Need help with linear algebra question!! Determine whether linear operator T:R^2 --> R^2 defined by the equation is one to one if so find the standard matrix for the inverse operator and find T^-1(w_1, w_2)

OpenStudy (anonymous):

Don't get how to find if it's one to one..

OpenStudy (anonymous):

What is T?

OpenStudy (anonymous):

Transformation? It's just showing that it's going from R^2 to R^2

OpenStudy (anonymous):

its says its defined by the equation....what equation?

OpenStudy (anonymous):

HAHA oh snap yeahh

OpenStudy (anonymous):

w_1 = 4x_1 - 6x_2 and w_2=2x_1+3x_2

OpenStudy (anonymous):

So we have: \[T(x_1,x_2) = (4x_1-6x_2, 2x_1+3x_2)\] To show that something is one-to-one, you need to prove that: \[T(a_1,a_2) = T(b_1,b_2) \Rightarrow (a_1,a_2) = (b_1,b_2)\]

OpenStudy (anonymous):

So lets start with the left hand side and see if we can produce the right hand side: \[T(a_1,a_2) = T(b_1,b_2)\] Using the definition of the transformation (the formula): \[T(a_1,a_2) = T(b_1,b_2) \Rightarrow (4a_1-6a_2, 2a_1+3a_2) = (4b_1-6b_2, 2b_1+3b_2)\]

OpenStudy (anonymous):

Sorry I can't see the formula so it's T(a_1, a_2) = T(b_1, b_2) --->(4a_1-6a_2, 2a_1+3a_2) = (4b_1-6b_2, 2b_1+3b_2)?

OpenStudy (anonymous):

Right. Am i doing something wrong that you cant see the formula?

OpenStudy (anonymous):

no it's just a lot of brackets and it says rightarrow haha but it's alright. Okay so how do you know a = b?

OpenStudy (anonymous):

Next, lets subtract both sides of that equation by the right hand side: \[(4a_1-6a_2, 2a_1+3a_2) = (4b_1-6b_2, 2b_1+3b_2) \Rightarrow \] \[(4a_1-6a_2, 2a_1+3a_2) - (4b_1-6b_2, 2b_1+3b_2) = 0\] combining them we get: \[((4a_1-6a_2)-(4b_1-6b_2), (2a_1+3a_2)-(2b_1+3b_2)) = (0, 0)\]

OpenStudy (anonymous):

Basically what you want to do is move things around, get a bunch of terms like (a_1 - b_1), (a_2-b_2), etc, and show that because the right hand side is 0, each of those terms must be 0. You end up with each a_i-b_i = 0 which is the same as a_1 = b_i

OpenStudy (anonymous):

i'll write the rest out on paper, one sec.

OpenStudy (anonymous):

can you just combine the a_1 and b_1 like that?

OpenStudy (anonymous):

like what?

OpenStudy (anonymous):

Like it was originally this right (4a_1-6a_2, 2a_1+3a_2) - (4b_1-6b_2, 2b_1+3b_2) = 0

OpenStudy (anonymous):

then you switch 2a_1 +3a_2 with 4b..etc to get (4a_1-6a_2)-(4b_1-6b_2), (2a_1+3a_2)-(2b_1+3b_2)) = (0, 0)

OpenStudy (anonymous):

yes, you could have combined them like that. my professors yell at me if i skip steps so ive gotten accustomed to showing everything lol

OpenStudy (anonymous):

Oh okay didn't know you could, no your steps really do help and sorry but do also know what it means when it says T^-1(w_1, w_2)?

OpenStudy (anonymous):

scanning a pic, one sec

OpenStudy (anonymous):

Oh okay I got up to finding the inverse of A then didnt know what to do.. thanks so much

OpenStudy (anonymous):

Hey joemath.. do you think I could trouble you with one more question you seem to know your linear algebra quite well

OpenStudy (anonymous):

Go ahead, shoot away :)

OpenStudy (anonymous):

okay this has to with basis and dimensions find the canonical basis for the solution space of the homogeneous system and state the dimension of the space..

OpenStudy (anonymous):

x_1 + x_2 + x_3 =0; 2x_1-x_2+4x_3 =0; 3x_1 + x_2 + 11x+3 =0

OpenStudy (anonymous):

I don't understand dimension of space

OpenStudy (anonymous):

Alright, im going to post something in a sec. Basically what you want to do is create the matrix that represents those equations, row reduce it, and look for the pivot and free variables.

OpenStudy (anonymous):

Ignore whats in the top left and top right.

OpenStudy (anonymous):

Haha okay.. I wrote the equation down wrong but it's alright

OpenStudy (anonymous):

so after you get the equations

OpenStudy (anonymous):

you put it into a vector form?

OpenStudy (anonymous):

yes. make the matrix that would represent that system. Then row reduce it.

OpenStudy (anonymous):

oh okay I think I got it thanks!

OpenStudy (anonymous):

no prob :)